OpenStudy (anonymous):

The length of a rectangle is three more than twice its width. The perimeter is 36 inches. Find the length and width. w = 5, l = 13 w = 8, l = 10 w = 6, l = 12 w = 7, l = 11 @iGreen

OpenStudy (shinalcantara):

the perimeter of a rectangle is P = 2L + 2W since the length is 3 more than twice the width you'll have: L = 2W + 3 figure it out

OpenStudy (shinalcantara):

you're given with the perimeter of 36 inches

OpenStudy (igreen):

Which option has the length as 3 more than twice its width"?

OpenStudy (igreen):

All of them have a Perimeter of 36 I think.

OpenStudy (igreen):

The answer is obviously A.. Width = 5 Length = 13 Length = Width * 2 + 3 13 = 5 * 2 + 3 13 = 10 + 3 13 = 13 And: P = 2(5) + 2(13) P = 10 + 26 P = 36

OpenStudy (igreen):

@DUDE..LOOK..ITS..SGTC